OpenStudy (anonymous):

two masses m(A) and m(B) moving with velocities V(A) and V(B) in opposite directions collide elastically and after that m(A) and m(B) with velocities V(B) and V(A) respectively. then the ratio m(A)/m(B) is ??

OpenStudy (anonymous):

while solving plz also explain me this ques too i dont get this completely

OpenStudy (amistre64):

hmm, ill have to look up what it means to collide elastically; i think its going to be they merge into one tho ....

OpenStudy (anonymous):

ok u need ans as i have an key acc. to that u can work ans:1

OpenStudy (amistre64):

http://www.launc.tased.edu.au/online/sciences/physics/elas.htm this helps me alittle since i like to play pool :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i be back in 5min ok

OpenStudy (amistre64):

class is starting for me so hopefully this gets resolved in the next 2 hours or so :) good luck tho the ratio of masses to me doesnt change if im seeing this right; or does it refer to the ration of the velcoities? either way, i gotta head out

OpenStudy (anonymous):

ok np actaully today i had my xams and dese are some que which i m unable to solve np thnx for coming :)

OpenStudy (mani_jha):

m(A)v(A)+m(B)v(B)=m(A)v(B)+m(B)v(A) m(A)(v(a)-v(b))=m(b)(v(a)-v(b)) mA=mB

OpenStudy (mani_jha):

For your exam, remember that after an elastic collision between two equal masses, the velocities get exchanged. For example, if a moving body hits another body which is at rest(of equal mass), then after the collison, the moving body comes to rest and it gives all of its energy to the body at rest, which starts moving with the same velocity. Thus, kinetic energy is conserved.

OpenStudy (anonymous):

ok and thanx..